OpenStudy (anonymous):

If v1=(2,-5) and v2=(4,-3), then the angle between the vectors is _____. Round your answer to two decimal places.

OpenStudy (anonymous):

\[v_1\cdot v_2=|v_1|~|v_2|~\cos\theta\]

OpenStudy (anonymous):

Do I add any of terms together or something?

OpenStudy (anonymous):

You could say that. Do you know how to find the dot product of two vectors?

OpenStudy (anonymous):

Like 2 + 4? and -5 + -3?

OpenStudy (anonymous):

Not quite, no. Given \(a=\langle a_1,a_2\rangle\) and \(b=\langle b_1,b_2\rangle\), then the dot product of \(a\) and \(b\) is \(a\cdot b=a_1b_1+a_2b_2\)

OpenStudy (anonymous):

so the answer is 23?

OpenStudy (psymon):

Yep

OpenStudy (psymon):

Now just get ||u|| and ||v||

OpenStudy (anonymous):

uh how do I do that?

OpenStudy (psymon):

It's pythagorean theorem actually. \[||u||=\sqrt{(a)^{2}+(b)^{2}}\] whichever letters you like. It's just like a^2 + b^2 = c^2. So for u, calculate: \[||u||= \sqrt{(2)^{2}+(-5)^{2}}\]

OpenStudy (anonymous):

i got 5.38

OpenStudy (anonymous):

and for v i got 5

OpenStudy (psymon):

So multiply those two results and then divide that into 23 and that'll = cos(theta)

OpenStudy (anonymous):

is it .39?

OpenStudy (anonymous):

wait was I suppose to take the answer from dividing it by 23 and put that answer into cos(theta)? @Psymon

OpenStudy (psymon):

\[\frac{ 23 }{ (5)(5.38) }=\cos \theta \]

OpenStudy (anonymous):

24.75

OpenStudy (psymon):

For theta?

OpenStudy (psymon):

I got 31.24.

OpenStudy (anonymous):

for 23/(5)95.38)

OpenStudy (psymon):

23 is divided by the product of those 2.

OpenStudy (anonymous):

Oh so 31.24 would be my final answer?

OpenStudy (psymon):

And again, I think it's a good habit to keep it in radical form as long as possible xD And yes, it would.

OpenStudy (anonymous):

thanks!

OpenStudy (psymon):

Yep yep.